Betty's Diner makes 15% profit in its lunches and 22% profit on its dinners. If the diner took in 2700 dollars and made 524 dollars profit, how much was spent at lunch. Let x represent the sales at lunch and y represent the sales at dinner. Write two equations, then solve.
Answer by quentin252(19)

x+y=2700 total sales
15% of x + 22% of y = .15x + .22y = 524 profits
so...from 1st eq
y=2700-x
plug into 2nd
.15x + .22(2700-x) = 524
.15x-.22x+594 = 524
-.07x = -70
x = 1000